Investments. A self employed contractor nearing retirement made 2 investments totaling $15,000. In 1 year these investments yielded $1023 in simple interest. Part of the money was invested at 6% and the rest at 7.5%. How much was invested at each rate ?
Answer by nerdybill(7384)

Investments. A self employed contractor nearing retirement made 2 investments totaling $15,000. In 1 year these investments yielded $1023 in simple interest. Part of the money was invested at 6% and the rest at 7.5%. How much was invested at each rate ?
.
Let x = amount invested at 6%
then
15000-x = amount invested at 7.5%
.
.06x + .075(15000-x) = 1023
.06x + 1125 - .075x = 1023
1125 - .015x = 1023
-.015x = -102
x = $6800 (amount invested at 6%)
.
Amount invested at 7.5%:
15000-x = 15000-6800 = $8200